# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ 3.3 FJ optimality condition for abstract problem (26)
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$ (42)
$\quad-\left.\sum_{i\in I_{1}}\int_{0}^{T}\left\langle\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s),\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}t-\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right].$
where we have used the Fubini theorem and the fact that the integral $\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)$ can be viewed as a Bochner integral valued in $\mathbb{R}^{l}$. Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$. Then, the arbitrariness of $\alpha\in\mathscr{U}[0,T]$ in (3.3) yields that, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
@@
$\qquad-\left.\sum_{i\in I_{1}}\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (43)
Using the representation (35) of the Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$, we obtain the stochastic minimum condition (3.11) from (3.3) by noting that $\lambda^{1}=Y$ and $\lambda^{2}=Z$ where the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ satisfies the BSDE (3.11).
On the other hand, the Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for $i\in I_{1}$, which is a consequence of the first complementary slackness condition given in (33) of Lemma 3.10. Moreover, $\eta_{t}^{j}$ vanishes on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))\neq 0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$, which results from the second complementary slackness condition in (33) of Lemma 3.10. In addition, we have that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$ as $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$. Finally, the uniqueness of solution to BSDE (3.11) is a result of the classical BSDE theory (c.f. Pham [37]). ∎
@@
According to the proof of Theorem 3.11, the adjoint process $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ in the constrained BSDE (3.11) can be actually interpreted as the components of the multiplier with respect to the McKean-Vlasov SDE (equality) constraint.
Remark 3.14.
In fact, Theorem 3.11 can be also applied to C-(E)MFC problem with Poisson jumps, which is formulated as follows:
$\begin{cases}\text{{\rm minimize} $J(\alpha)$ {\rm over} $\alpha\in\mathscr{U}[0,T];$}\\[6.00006pt] \displaystyle\text{{\rm s.t.} $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}W_{s}$}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$+\int_{0}^{t}\int_{Z}\gamma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s},z)\tilde{N}(\mathrm{d}s,\mathrm{d}z)$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-{\rm a.s}.}\end{cases}$ (44)
Here, $\tilde{N}(\mathrm{d}s,\mathrm{d}z):=N(\mathrm{d}s,\mathrm{d}z)-\lambda_{Z}(\mathrm{d}z)\mathrm{d}s$ is a compensated Poisson random measure on some measurable space $(Z,\mathscr{Z})$ with finite intensity measure $\lambda_{Z}(\mathrm{d}z)\mathrm{d}s$, and the jump coefficient $\gamma:[0,T]\times\mathbb{R}^{n}\times{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times Z\mapsto\mathbb{R}^{n}$ is assumed to satisfy
  • (B1)
    $\gamma$ is borel measurable and continuously differentiable w.r.t. $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Furthermore, $\gamma$ admits a continuous L-derivative w.r.t. the measure $\rho\in{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$.
  • (B2)
    there is a constant $M>0$ such that, for all $(x,\rho),(x^{\prime},\rho^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ and $(t,u,z)\in[0,T]\times\mathbb{R}^{l}\times Z$,
    $\left|\gamma(t,x^{\prime},\rho^{\prime},u,z)-\gamma(t,x,\rho,u,z)\right|$ $\leq M\left(|x-x^{\prime}|+\mathcal{W}_{2}(\rho,\rho^{\prime})\right),$
    $\left|\gamma(t,x,\rho,u,z)\right|+\left|\nabla_{u}\gamma(t,x,\rho,u,z)\right|$ $\leq M\left(1+|x|+|u|+M_{2}(\rho)\right).$
Correspondingly, the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ is defined as the natural filtration generated by the Brownian motion $W$ and the Poisson random measure $N(\mathrm{d}s,\mathrm{d}z)$. Then, the space $\mathbb{M}^{2}$ corresponding to the McKean-Vlasov jump diffusion state process in (44) should be modified to the set of $\mathbb{F}$-adapted (right continuous with left limits) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ admitting the following Itô’s representation:
$\eta_{t}=\eta_{0}+\int_{0}^{t}\eta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\eta_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{t}\int_{Z}\eta_{s}^{3}(z)\tilde{N}(\mathrm{d}s,\mathrm{d}z),$ (45)
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $\eta^{1}=(\eta_{t}^{1})_{t\in[0,T]}$ and $\eta^{2}=(\eta_{t}^{2})_{t\in[0,T]}$ taking values respectively on $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$ are $\mathbb{F}$-adapted processes such that $\mathbb{E}[\int_{0}^{T}(|\eta_{t}^{1}|^{2}+|\eta_{t}^{2}|^{2})\mathrm{d}t]<\infty$. Furthermore, the $\mathbb{R}^{n}$-valued process $\eta^{3}=(\eta_{t}^{3}(z))_{(t,z)\in[0,T]\times Z}$ is $\mathbb{F}$-adapted and satisfies that $\mathbb{E}[\int_{0}^{T}\int_{Z}|\eta^{3}_{t}(z)|^{2}\lambda_{Z}(\mathrm{d}z)\mathrm{d}t]<\infty$. Thus, $\mathbb{M}^{2}$ becomes a Hilbert space under the modified inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by, for all $\eta,\zeta\in\mathbb{M}^{2}$,
$\langle\eta,\zeta\rangle_{\mathbb{M}^{2}}$ $:=\mathbb{E}\left[\langle\eta_{0},\zeta_{0}\rangle+\int_{0}^{T}\left(\langle\eta_{t}^{1},\zeta_{t}^{1}\rangle+\mathrm{tr}\left(\eta_{t}^{2}(\zeta_{t}^{2})^{\top}\right)\right)\mathrm{d}t+\int_{0}^{T}\int_{Z}\langle\eta^{3}_{t}(z),\zeta^{3}_{t}(z)\rangle\lambda_{Z}(\mathrm{d}z)\mathrm{d}t\right].$ (46)
Note that Lemma 3.5 still holds because $\mathcal{L}(X_{t})$ is continuous (c.f. [9]). Moreover, Lemma 3.8 and Lemma 3.9 in this jump-diffusion framework are also valid. One can deduce the SMP for problem (44) by following the proof of Theorem 3.11: there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that conditions (i)-(iii) of Theorem 3.11 hold and the stochastic (first-order) minimum condition is given by, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t},p_{t},P_{t},K_{t})+\mathbb{E}^{\prime}\left[\partial_{\nu}H^{r_{0}}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime},p_{t},P_{t},K_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\right.$
$\qquad\left.-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (47)
The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\times L^{2}((Z,\mathscr{Z},\lambda_{Z});\mathbb{R}^{n})\mapsto\mathbb{R}$ is defined by
$H^{r_{0}}(t,x,\rho,u,p,P,K)$ $:=\langle b(t,x,\rho,u),y\rangle+\mathrm{tr}\left(\sigma(t,x,\rho,u)z^{\top}\right)+\int_{Z}\gamma(t,x,\rho,u,z)K(z)\lambda_{Z}(\mathrm{d}z)$
$\quad+r_{0}f(t,x,\rho,u),$ (48)
and the process triplet $(p,P,K)=(p_{t},P_{t},K_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\times L^{2}((Z,\mathscr{Z},\lambda_{Z});\mathbb{R}^{n})$ is the unique solution to the following BSDE:
$p_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime})\right]+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s},p_{s},P_{s},K_{s})\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime},p_{s},P_{s},K_{s})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}-\int_{t}^{T}\int_{Z}K_{s}(z)\tilde{N}(\mathrm{d}s,\mathrm{d}z).$ (49)
In particular, when $m=0$, similar to Remark 3.12, the above BSDE (3.14) reduces to Eq. (36) in [9] in the McKean-Vlasov jump-diffusion model.
@@ 4.1 Constrained stochastic control problem
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s..}\end{cases}$ (50)
@@
    the positive Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (53)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\int_{[t,T]}\nabla_{x}\phi(s,\widehat{X}_{s})\mu^{i}(\mathrm{d}s)$
$\quad-\int_{t}^{T}Z_{s}\mathrm{d}W_{s},$ (57)
@@
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$X_{t}^{\alpha}\geq a_{t}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (58)
@@
where $A_{t}=\int_{0}^{t}\eta_{s}\mathrm{d}s$ is the adapted increasing process with $\eta=(\eta_{t})_{t\in[0,T]}$ being the multiplier given in Corollary 4.1 such that
$\mathbb{E}\left[\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}\right]=0.$
This also implies that, $\mathbb{P}$-$\mathrm{a.s.}$,
$\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}=0,$ (60)
in view of $\widehat{X}_{t}-a_{t}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.. Therefore, we can see that Eq. (59) resembles a reflected BSDE with the reflecting barrier $t\mapsto a_{t}$.

### v2

@@ 3.3 FJ optimality condition for abstract problem (19)
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$ (34)
$\quad-\sum_{i\in I_{1}}\int_{0}^{T}\left\langle\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s),\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}t$
$\quad-\left.\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\left(\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\nu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\right)\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right],$
where we used the Fubini theorem and the fact that $\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)$ can be viewed as a Bochner integral valued in $\mathbb{R}^{l}$. Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$. Then, the arbitrariness of $\alpha\in\mathscr{U}[0,T]$ in (3.3) yields that
@@
$\qquad-\left.\sum_{i\in I_{1}}\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}\right.$
$\qquad-\left.\sum_{j\in J_{2}}\mathbb{E}^{\prime}\left[\partial_{\nu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,~{}\forall u\in U,\quad\mathrm{d}t\times\mathrm{d}\mathbb{P}\mbox{-}\mathrm{a.s.}$ (35)
Using the representation (27) of the Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$, we obtain the stochastic minimum condition (3.12) from (3.3) by noting that $\lambda^{1}=Y$ and $\lambda^{2}=Z$, where the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ satisfies the BSDE (3.12).
On the other hand, the Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for $i\in I_{1}$, which is a consequence of the first complementary slackness condition given in (25) of Lemma 3.11. Moreover, $\eta_{t}^{j}$ vanishes on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\rho}_{t},\widehat{\alpha}_{t}(\omega))\neq 0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$, which results from the second complementary slackness condition in (25) of Lemma 3.11. In addition, we have that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$ as $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$. Finally, the uniqueness of solution to BSDE (3.12) is a result of the classical BSDE theory (c.f. Pham [38]). ∎
@@
According to the proof of Theorem 3.12, the adjoint process $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ in the constrained BSDE (3.12) can be interpreted as the components of the Lagrange multiplier with respect to the McKean-Vlasov SDE (equality) constraint.
@@ 4.1 Constrained stochastic control problem
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[3.99994pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[6.00006pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s..}\end{cases}$ (36)
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (39)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\int_{[t,T]}\nabla_{x}\phi(s,\widehat{X}_{s})\mu^{i}(\mathrm{d}s)-\int_{t}^{T}Z_{s}\mathrm{d}W_{s},$
@@
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[3.99994pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$X_{t}^{\alpha}\geq a_{t}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (43)
@@
where $A_{t}=\int_{0}^{t}\eta_{s}\mathrm{d}s$ is the adapted increasing process with $\eta=(\eta_{t})_{t\in[0,T]}$ being the multiplier given in Corollary 4.1 such that $\mathbb{E}[\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}]=0$. This also implies that, $\mathbb{P}$-$\mathrm{a.s.}$, $\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}=0$ in view of $\widehat{X}_{t}-a_{t}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.. Therefore, we can see that Eq. (44) resembles a reflected BSDE with the reflecting barrier $t\mapsto a_{t}$.